OpenStudy (vshiroky):

Need answer verification please (2y)^2+4y When y=-3 We got 24

OpenStudy (usukidoll):

when y = -3 we have \[(2(-3))^2+4(-3)\] \[(-6)^2 -12 \] \[36-12 = 24\]

OpenStudy (vshiroky):

Could you check another?

OpenStudy (usukidoll):

sure

OpenStudy (vshiroky):

Thank you, I'm trying to help her finish hers so I can start on mine lol

OpenStudy (vshiroky):

m^3+3(x+2) when m=-2 and x=4

OpenStudy (vshiroky):

I got 10, she got 14

OpenStudy (usukidoll):

ok so first we need to plug in m=-2 and x = 4 \[m^3+3(x+2) \] \[(-2)^3+3(4+2)\] \[\[-8+3(6)\]\] \[-8+18 = 10 \]

OpenStudy (vshiroky):

Do you care if I keep typing more?

OpenStudy (usukidoll):

depends on the topic and difficulty how many more problems?

OpenStudy (vshiroky):

Same stuff and I will take whatever help you can give until you dont want to anymore because there is quite a few more

OpenStudy (usukidoll):

ok...

OpenStudy (vshiroky):

6k^3+3d k=-1 and d=-2 We got -12

OpenStudy (usukidoll):

\[6k^3+3d \] for k=-1 and d = -2 \[6(-1)^3+3(-2) \] \[6(-1)+3(-2) \rightarrow -6-6 =-12\]

OpenStudy (vshiroky):

Thanks... (4+x)^2/x-6 when x=6 We got 100

OpenStudy (usukidoll):

\[\frac{(4+x)^2}{x-6}\] for x = 6.. wait a minute... this is undefined... if we let x = 6 in the denominator we will have 100/0 and we can't have a 0 in the denominator.

OpenStudy (vshiroky):

oops it was supposed to be x-1 on the bottom

OpenStudy (vshiroky):

so 20 is the answer

OpenStudy (usukidoll):

oH! xxxxxxxxxxxxxxD

OpenStudy (usukidoll):

\[\frac{(4+x)^2}{x-1}\] for x = 6 100/5 = 20 yup !

OpenStudy (vshiroky):

The next one is 4+7/x-2 + 3 when x=5 We got 6 2/3

OpenStudy (usukidoll):

is it \[4 + \frac{7}{x}-2 +3\] or \[4 + \frac{7}{x-2} +3\] ???

OpenStudy (usukidoll):

I don't understand the format of the question... could you draw it out?

OpenStudy (vshiroky):

Ill just skip that one I cant figure out how to type it lol

OpenStudy (vshiroky):

I'll try

OpenStudy (vshiroky):